what is x in the equation 5 times 7x = 1

Found 2 solutions by checkley79, Alan3354:
Answer by checkley79(3341)

5*7X=1
35X=1
X=1/35 ANS.

Answer by Alan3354(69443)

what is x in the equation 5 times 7x = 1
---------
5*7x = 1
35x = 1
x = 1/35